OpenStudy (anonymous):

The ratio of the radii of two spheres is 2:5. The volume of the smaller sphere is 192 in3. What is the volume of the larger sphere?

OpenStudy (anonymous):

Ratio of volumes is 2^3:5^3 or 8:125 192/V = 8/125 V = 3000 in^3